OpenStudy (anonimous):

Jack has a fair coin. He tosses the coin four times. What is the probability that Jack gets heads, then heads, then tails, and then tails? In fraction

Vocaloid (vocaloid):

what's the probability of getting heads?

OpenStudy (anonimous):

idk :)

Vocaloid (vocaloid):

how many sides on a coin?

OpenStudy (anonimous):

2

Vocaloid (vocaloid):

good, and how many sides are heads?

OpenStudy (anonimous):

1

Vocaloid (vocaloid):

good! so our probability of getting one heads is 1/2

Vocaloid (vocaloid):

what about tails?

OpenStudy (anonimous):

idk, may be 1

Vocaloid (vocaloid):

remember, how many sides?

OpenStudy (anonimous):

2

Vocaloid (vocaloid):

good, and 1 of them is tails

Vocaloid (vocaloid):

so 1/2

Vocaloid (vocaloid):

now we just multiply:

OpenStudy (anonimous):

ok, thank you sooo much I appreciate it so much :)

Vocaloid (vocaloid):

(heads * heads* tails * tails) = (1/2) * (1/2) * (1/2) * (1/2) = ?

OpenStudy (anonimous):

4?

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

what is 2 * 2 * 2 * 2 = ?

Vocaloid (vocaloid):

* means multiply

OpenStudy (anonimous):

6

Vocaloid (vocaloid):

almost

OpenStudy (anonimous):

I am so bad in multiplication!!!!!!!!!!!!

OpenStudy (anonimous):

12?

Vocaloid (vocaloid):

start with 2 double it double it again double it again

Vocaloid (vocaloid):

what is 2 * 2 = ?

OpenStudy (anonimous):

or 9?

OpenStudy (anonimous):

it's 9

Vocaloid (vocaloid):

not 9

Vocaloid (vocaloid):

what is 2 *2 ?

OpenStudy (anonimous):

4

Vocaloid (vocaloid):

good what is 4 * 2 = ?

OpenStudy (anonimous):

8

Vocaloid (vocaloid):

good, now what is 8 * 2 ?

OpenStudy (anonimous):

16

Vocaloid (vocaloid):

fantastic! so 2 * 2 * 2 * 2 = 16 which means (1/2) * 1/2 * 1/2 * 1/2 = 1/16 = your answer

OpenStudy (anonimous):

1/16 is the answer?

Vocaloid (vocaloid):

yes

OpenStudy (anonimous):

ok :)